# Overcoming label shift in targeted federated learning

### v1

@@ Abstract
Federated learning enables multiple actors to collaboratively train models without sharing private data. This unlocks the potential for scaling machine learning to diverse applications. Existing algorithms for this task are well-justified when clients and the intended target domain share the same distribution of features and labels, but this assumption is often violated in real-world scenarios. One common violation is label shift, where the label distributions differ across clients or between clients and the target domain, which can significantly degrade model performance. To address this problem, we propose FedPALS, a novel model aggregation scheme that adapts to label shifts by leveraging knowledge of the target label distribution at the central server. Our approach ensures unbiased updates under stochastic gradient descent, ensuring robust generalization across clients with diverse, label-shifted data. Extensive experiments on image classification demonstrate that FedPALS consistently outperforms standard baselines by aligning model aggregation with the target domain. Our findings reveal that conventional federated learning methods suffer severely in cases of extreme client sparsity, highlighting the critical need for target-aware aggregation. FedPALS offers a principled and practical solution to mitigate label distribution mismatch, ensuring models trained in federated settings can generalize effectively to label-shifted target domains.
@@ 1 Introduction
Federated learning has become a prominent paradigm in machine learning, enabling multiple clients to collaboratively train models without sharing their data (Kairouz et al., 2021). Central to this development is the widely-used federated averaging (FedAvg) algorithm, which aggregates model updates from clients, weighted by their data set sizes (McMahan et al., 2017). This aggregation rule is well-justified when client data is independent and identically distributed (i.i.d.) and has been effective in diverse domains such as healthcare (Sheller et al., 2020), finance risk prediction (Byrd and Polychroniadou, 2020), and natural language processing (Hilmkil et al., 2021).
The justification for FedAvg is weaker when clients exhibit systematic data heterogeneity, such as in the case of label shift (Zhao et al., 2018; Woodworth et al., 2020), since the learning objectives of clients differ from the objective optimized by the central server. Consider a federated learning task involving multiple retail stores (clients) where the goal is to predict the sales of different products based on customer purchase history, and deploy the trained model in a new store (target). Each store’s sales reflect local preferences, leading to differences in label distributions and empirical risks. When label distributions vary substantially, the performance of FedAvg can be severely hampered (Karimireddy et al., 2020; Zhao et al., 2018). So, how can models be trained effectively across such heterogeneous clients? Several methods have been proposed to deal with client heterogeneity, such as regularization (Li et al., 2020a, 2021), clustering (Ghosh et al., 2020; Vardhan et al., 2024), and meta-learning approaches (Chen et al., 2018; Jiang et al., 2019; Fallah et al., 2020a). Still, these studies and prior work in federated learning assume that the target (or test) domain shares the same distribution as the combined training data from the clients. In many real-world applications, models must generalize to new target domains with different distributions and without training data.
In the retail example, the target store has no individual transaction data available and is not involved in training the model. The challenge of generalizing under distributional domain shift without target data is not exclusive to federated learning but arises also in centralized settings (Blanchard et al., 2011; Muandet et al., 2013; Ganin et al., 2016). In centralized scenarios, domain adaptation techniques such as sample re-weighting (Lipton et al., 2018a), or domain-invariant representations (Arjovsky et al., 2020) can be employed to mitigate the effects of distributional shifts. However, these approaches require central access to observations from both the source and target domains. This is not feasible in federated learning due to the decentralized nature of the data. Neither the server nor the clients have access to both data sets, making direct application of these methods impractical.
@@
This work aims to improve the generalization of federated learning to target domains under label shift, in settings where the different label distributions of clients and target domains are known to the central server but unknown to the clients (see Section 2). To address this problem, we propose a novel aggregation scheme called FedPALS that optimizes a convex combination of client models to ensure that the aggregated model is better suited for the label distribution of the target domain (Section 3.1). Our approach is both well-justified and practical. We prove that the resulting stochastic gradient update behaves, in expectation, as centralized learning in the target domain (Proposition 1), and examine its relation to standard federated averaging (Proposition 3.2). We demonstrate the effectiveness of FedPALS through an extensive empirical evaluation (Section 5), showing that it outperforms traditional approaches in scenarios where distributional shifts pose significant challenges. Moreover, we observe that traditional methods struggle particularly in scenarios where the training clients have sparse label distributions. As we show in Figures 2(c) and 6(c) in Section 5, performance drops sharply when a majority of labels are completely unrepresented in clients, highlighting the limitations of existing approaches under label shift with extreme client sparsity.
2 Targeted federated learning under label shift
@@
Federated learning is a distributed machine learning paradigm wherein a global model $h_{\theta}$ is trained at a central server by aggregating parameter updates from multiple clients (McMahan et al., 2017). We focus on classification tasks in which the goal is for $h_{\theta}$ to predict the most probable label $Y\in\{1,...,K\}$ for a given input $X\in\mathcal{X}$. Each client $i=1,...,M$ holds a data set $D_{i}=\{(x_{i,1},y_{i,1}),...,(x_{i,n_{i}},y_{i,n_{i}})\}$ of $n_{i}$ labeled examples. Due to privacy concerns or limited communication bandwidth, these data sets cannot be shared with other clients or with the central server. Learning proceeds over rounds $t=1,...,t_{max}$, each completing three steps: (1) The central server broadcasts the current global model parameters $\theta_{t}$ to all clients; (2) Each client $i$ computes updated parameters $\theta_{i,t}$ based on their local data set $D_{i}$ and sends these updates back to the server; (3) The server aggregates the clients’ updates to obtain the new global model $\theta_{t+1}$.
Client samples $D_{i}$ are assumed to be drawn i.i.d. from a local distribution $S_{i}(X,Y)$. A common assumption in federated learning is that all client distributions are identical, i.e., $\forall i,j\in[M]:S_{i}=S_{j}$. Under this assumption, the learned central model is expected to perform well on average across clients. We can think of this as a special case of an intended target domain $T(X,Y)$ which coincides with the marginal distribution of clients, $\bar{S}=\sum_{i=1}^{M}\frac{n_{i}}{N}S_{i}$, where $N=\sum_{i=1}^{M}n_{i}$. But, in general, the target domain $T(X,Y)$ may be distinct from individual clients and their aggregate (Bai et al., 2024). Moreover, clients may exhibit significant heterogeneity in their distributions, $S_{i}\neq S_{j}$ (Karimireddy et al., 2020; Li et al., 2020b). We refer to this problem as targeted federated learning.
We study targeted federated learning under known label shift, where no samples from the joint distribution $T(X,Y)$ are available but the target label distribution $T(Y)$ is known to the server.
@@
Let $\mathcal{X}\subset\mathbb{R}^{d}$ denote the $d$-dimensional input space and $\mathcal{Y}=\{1,...,K\}$ the label space. Given a set of clients $S_{1},...,S_{M}$ and a target domain $T$, our objective is to minimize the expected risk, $R_{T}$ of a classifier $h_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, with respect to a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$ over the target distribution $T$,
[…]
Obtaining a good estimate of $T(Y)$ is often feasible since it represents aggregate statistics that can be collected without the need for a large dataset. In our retail example, $Y=y$ would represent a sale in a specific product category, and $T(y)$ would correspond to the proportion of total sales in that category. A company could estimate these proportions without logging customer features $X$. Our setting differs from domain generalization which lacks a specific target domain (Bai et al., 2024).
@@
In our setting, the target and all client label distributions are all assumed distinct: $\forall i\neq j\in[M]:S_{i}(Y)\neq S_{j}(Y)\neq T(Y)$. Furthermore, the target distribution is assumed to differ from the client aggregate, $T(Y)\neq\bar{S}(Y)$. While the server has access to all marginal label distributions ${S_{i}(Y)}_{i=1}^{M}$ and $T(Y)$, these are not available to the clients. For instance, most retailers would be hesitant to share their exact sales statistics $T(Y)$ with competitors. Instead, they might provide this information to a neutral third party (central server) responsible for coordinating the federated learning process. More broadly, clients $i\neq j$ do not communicate directly with each other but rather interact with the central server through model parameters. While it is technically possible for the server to infer each client’s label distribution $S_{i}(Y)$ based on their parameter updates (Ramakrishna and Dán, 2022), doing so would likely be considered a breach of trust in practical applications.
@@
The distributional shifts between clients and the target are restricted to label shift—while the label distributions vary across clients and the target, the class-conditional input distributions are identical.
@@
For the set of clients $S_{1},...,S_{M}$ and the target domain $T$,
@@
This setting has been well studied in non-federated learning, both in cases where the label marginals are known and where they are not (Lipton et al., 2018b). In the retail example, label shift would mean that while the proportion of sales across product categories ($S_{i}(Y)$ and $T(Y)$) varies between different retailers and the target, the purchasing patterns within each category ($S_{i}(X\mid Y)$ and $T(X\mid Y)$) remain consistent. In other words, although retailers may sell different quantities of products across categories, the characteristics of customers buying a particular product (conditional on the product category) are assumed to be the same.
Our central question is: In federated learning, how can we aggregate the parameter updates $\theta_{i,t}$ of the $M$ clients, whose data sets are drawn from distributions $S_{1},...,S_{M}$, such that the resulting federated learning algorithm minimizes the target risk, $R_{T}$? The fact that clients are ignorant of the shift in $T(Y)$ affects the optimal strategy; direct access to $T(Y)$ would allow sample re-weighting or upsampling in the client objectives (Rubinstein and Kroese, 2016). This is not possible here.
3 FedPALS: Parameter aggregation to adjust for label shift
In classical federated learning, all clients and the target domain are assumed i.i.d., and thus, the target risk ((1)) is equal to the expected risk in any client
$R_{T}(h)=\underset{(X,Y)\sim S_{i}}{\mathbb{E}}[\ell(h(X),Y)]\eqqcolon R_{i}(h),\;\;\mbox{for all}\;i=1,...,M\leavevmode\nobreak\ .$
[…]
As a result, if clients perform a single local gradient descent update, any convex combination of these gradients (updates) is equal in distribution to a classical (centralized) batch update for the target domain, given the previous parameter value,
[…]
where $\Delta^{M-1}=\{\alpha\in[0,1]^{M}:\sum_{i}\alpha_{i}=1\}$ is the simplex over $M$ elements. This property justifies the federated stochastic gradient (FedSGD) and FedAvg principles (McMahan et al., 2017), [^1] Strictly speaking, only FedSGD uses the property directly, but FedAvg is a natural extension. which aggregate model parameter updates through a convex combination, chosen to give weight to clients proportional to their sample size,
@@ 3 FedPALS: Parameter aggregation to adjust for label shift
A limitation of FedAvg aggregation is that when client and target domains are not identically distributed, and $T\neq\bar{S}$, $\{\nabla\hat{R}_{i}\}_{i=1}^{M}$ are no longer unbiased estimates of the risk gradient in the target domain, $\nabla R_{T}$. As a result, the FedAvg update is not an unbiased estimate of a model update computed in the target domain. As we see in Table 1, this can have large effects on model quality.
3.1 Overcoming label shift while controlling variance
Next, we develop an alternative aggregation strategy that partially overcomes the limitations of FedAvg under label shift (Assumption 1). Under label shift, it holds that
[…]
In centralized (non-federated) learning under label shift, this insight is often used to re-weight (Lipton et al., 2018b) or re-sample (Japkowicz and Stephen, 2002) the empirical risk in the source domain. This is not an option here since $T(Y)$ is not revealed to the clients. For now, assume instead that the target label distribution can be covered by a convex combination of client label distributions. For this, let $\mathrm{Conv}(S)$ denote the convex hull of distributions $\{S_{i}(Y)\}_{i=1}^{M}$.
@@ 3.1 Overcoming label shift while controlling variance
For the label marginals $S_{1}(Y),...,S_{M}(Y),T(Y)\in\Delta^{K-1}$, the target label distribution $T(Y)$ is covered by a convex combination $\alpha^{c}$ of client label distributions,
$T\in\mathrm{Conv}(S),\quad\mbox{i.e.,}\quad\exists\alpha^{c}\in\Delta^{M-1}:T(y)=\sum_{i=1}^{M}\alpha_{i}^{c}S_{i}(y)\quad\forall y\in[K]\leavevmode\nobreak\ .$ (4)
Note that, under label shift, Assumption 2 implies that $T(X,Y)=\sum_{i=1}^{M}\alpha^{c}_{i}S_{i}(X,Y)$, as well. Thus, under Assumptions 12, we have for any $\alpha^{c}$ satisfying (4),
$R_{T}(h)=\sum_{y=1}^{K}\left(\sum_{i=1}^{M}\alpha^{c}_{i}S_{i}(y)\right)\mathbb{E}[\ell(h(X),y)\mid Y=y]=\sum_{i=1}^{M}\alpha^{c}_{i}R_{S_{i}}(h).$ (5)
By extension, aggregating client updates with weights $\alpha^{c}$ will be an unbiased estimate of the update.
@@
where $\theta_{t+1}^{T}=\theta_{t}-\eta\nabla_{\theta}\hat{R}_{T}(h_{\theta_{t}})$ is the batch stochastic gradient update for $\hat{R}_{T}$ that would be obtained with a sample from the target domain. A proof is given in Appendix C.
By Proposition 1, we are justified [^2] Note that, just like in FedAvg, aggregating client parameter updates resulting from multiple local stochastic gradient steps (e.g., one epoch) is not guaranteed to locally minimize the target risk. in replacing the aggregation step of FedAvg with one where clients are weighted by $\alpha^{c}$. However, Assumption 2 may not hold, and $\alpha^{c}$ may not exist. For instance, if the target marginal is sparse, only clients with the exact same sparsity pattern as $T$ can be used in a convex combination $\alpha^{c}S=T$. That is, if we aim to classify images of animals and $T$ contains no tigers, then no clients contributing to the combination can have data containing tigers. At least, since $\{S_{i}(Y)\}_{i=1}^{M},T(Y)$ are known to the server, it is straightforward to verify Assumption 2.
A pragmatic choice when Assumption 2 is violated is to look for the convex combination $\alpha^{0}$ that most closely aligns with the target label distribution, and use that in the aggregation step,
$\alpha^{0}=\operatorname*{arg\,min}_{\alpha\in\Delta^{M-1}}\bigg{\|}\sum_{i=1}^{M}\alpha_{i}S_{i}(Y)-T(Y)\bigg{\|}_{2}^{2}\quad\mbox{and}\quad\theta^{0}_{t+1}=\sum_{i}\alpha^{0}_{i}\theta_{i,t}\leavevmode\nobreak\ .$ (6)
@@
Figure 1: Illustration of the target label marginal $T$ and client marginals $S_{1},...,S_{4}$ in a ternary classification task, $Y\in\{0,1,2\}$. Left: there are fewer clients than labels, $M<K$, and $T\not\in\mathrm{Conv}(S)$; $\alpha^{0}S$ is a projection of $T$ onto $\mathrm{Conv}(S)$. Right: $T\in\mathrm{Conv}(S)$ and coincides with $\alpha^{0}S$. In both cases, the label marginal $\alpha^{FA}S$ implied by FedAvg is further from the target distribution.
@@
In light of the results above, we propose a client aggregation step such that the weighting of clients’ label distributions will a) closely align with the target label distribution, and b) minimize the variance due to weighting using the inverse of the ESS. For a given regularization parameter $\lambda\in[0,\infty)$, we define a weighting scheme $\alpha^{\lambda}$ as the solution to the following problem.
@@
and aggregate client parameters inside federated learning as $\theta^{\lambda}_{t+1}=\sum_{i=1}^{M}\alpha^{\lambda}_{i}\theta_{i,t}$. We refer to this strategy as Federated learning with Parameter Aggregation for Label Shift (FedPALS).

### v2

# Overcoming label shift with target-aware federated learning

@@ Abstract
Federated learning enables multiple actors to collaboratively train models without sharing private data. Existing algorithms are successful and well-justified in this task when the intended target domain, where the trained model will be used, shares data distribution with the aggregate of clients, but this is often violated in practice. A common reason is label shift—that the label distributions differ between clients and the target domain. We demonstrate empirically that this can significantly degrade performance. To address this problem, we propose FedPALS, a principled and practical model aggregation scheme that adapts to label shifts to improve performance in the target domain by leveraging knowledge of label distributions at the central server. Our approach ensures unbiased updates under federated stochastic gradient descent which yields robust generalization across clients with diverse, label-shifted data. Extensive experiments on image classification tasks demonstrate that FedPALS consistently outperforms baselines by aligning model aggregation with the target domain. Our findings reveal that conventional federated learning methods suffer severely in cases of extreme label sparsity on clients, highlighting the critical need for target-aware aggregation as offered by FedPALS.
@@ 1 Introduction
Federated learning (FL) has emerged as a powerful paradigm for training machine learning models collaboratively across multiple clients without sharing data (McMahan et al., 2017; Kairouz et al., 2021). This is attractive in problems where privacy is paramount, such as healthcare (Sheller et al., 2020), finance (Byrd and Polychroniadou, 2020), and natural language processing (Hilmkil et al., 2021). While effective when data from different clients are identically distributed, the performance of federated learning can degrade significantly when clients exhibit systematic data heterogeneity, such as label shift (Zhao et al., 2018; Woodworth et al., 2020).
Most federated learning research, even that addressing data heterogeneity, focuses on what we term standard federated learning, where the test distribution matches the combined distribution of training clients. However, many real-world applications require generalization to a target client or domain with a distinct, unknown data distribution. Consider a retail scenario: multiple stores (clients) collaboratively train a sales prediction model using their local purchase histories. While each store’s data reflects its unique customer base, the goal is to deploy the model in a new store (target client) with different customer preferences, and no historical records. This is target-aware federated learning, a more challenging paradigm than standard federated learning due to the inherent distributional shift between training and test data.
The problem of generalizing under distributional shifts has been extensively studied in centralized settings, often under the umbrella of domain adaptation (Blanchard et al., 2011; Ganin et al., 2016). However, traditional domain adaptation techniques, such as sample re-weighting (Lipton et al., 2018) or domain-invariant representation learning (Arjovsky et al., 2020), typically require access to data from both source and target domains. This requirement is incompatible with the decentralized nature of federated learning, where neither the server nor the clients share data between them. While several techniques have been proposed to address client heterogeneity in standard FL, such as regularization (Li et al., 2020a, 2021), clustering (Ghosh et al., 2020; Vardhan et al., 2024), and meta-learning (Chen et al., 2018; Jiang et al., 2019), they do not address the challenge of generalizing to a target client with differing data distribution.
@@
We introduce the problem of target-aware federated learning under label shift, where the goal is to train a model that generalizes well to a target client (domain) whose label distribution differs from those of the training clients (see Section 2). To address this problem, we propose a novel aggregation scheme called FedPALS that optimizes a convex combination of client models to ensure that the aggregated model is better suited for the label distribution of the target domain (Section 3). We prove that the resulting stochastic gradient update behaves, in expectation, as centralized learning in the target domain (Proposition 1), and examine its relation to standard federated averaging (Proposition 3.1). We demonstrate the effectiveness of FedPALS through an extensive empirical evaluation (Section 5), showing that it outperforms traditional approaches in scenarios where distributional shifts pose significant challenges, at the small cost of sharing client label marginals with the central server. Moreover, we observe that traditional methods struggle particularly in scenarios where training clients have sparse label distributions, highlighting the need for target-aware aggregation strategies.
2 Target-aware federated learning with label shift
@@
In federated learning, a global model $h_{\theta}$ is produced by a central server by aggregating updates to model parameters $\theta$ from multiple clients (McMahan et al., 2017). We focus on classification tasks in which the goal is for $h_{\theta}$ to predict the most probable label $Y\in\{1,...,K\}$ for a given $d$-dimensional input $X\in\mathcal{X}\subset\mathbb{R}^{d}$. Each client i=1,,Mi=1,...,M holds a data set $D_{i}=\{(x_{i,1},y_{i,1}),...,(x_{i,n_{i}},y_{i,n_{i}})\}$ of $n_{i}$ labeled examples, assumed to be drawn i.i.d. from a local client-specific distribution $S_{i}(X,Y)$. Due to constraints on privacy or communication, these data sets cannot be shared directly with other clients or with the central server.
Learning proceeds over rounds $t=1,...,t_{max}$, each comprising three steps: (1) The central server broadcasts the current global model parameters $\theta_{t}$ to all clients; (2) Each client $i$ computes updated parameters $\theta_{i,t}$ based on their local data set $D_{i}$, and sends these updates back to the server; (3) The server aggregates the clients’ updates, for example, using federated averaging (FedAvg) (McMahan et al., 2017) or related techniques, to produce the new global model $\theta_{t+1}$.
A common implicit assumption in federated learning is that the learned model will be applied in a target domain $T(X,Y)$ that coincides with the aggregate distribution of clients,
$\bar{S}(X,Y)=\sum_{i=1}^{M}\frac{n_{i}}{N}S_{i}(X,Y),$ (1)
where $N=\sum_{i=1}^{M}n_{i}$. To this end, trained models are evaluated in terms of their average performance over clients. However, in applications, the intended target domain may be different entirely (Bai et al., 2024). Here, we assume that the target domain is distinct from all client distributions: $\forall i:T(X,Y)\neq S_{i}(X,Y)$ and from the client aggregate, $T(X,Y)\neq\bar{S}(X,Y)$. We refer to this setting as target-aware federated learning. While distributional shift between clients is a well-recognized problem in federated learning, the target domain is still typically the client aggregate $\bar{S}$ (Karimireddy et al., 2020; Li et al., 2020b). Our setting differs also from federated domain generalization which lacks a specific target domain (Bai et al., 2024).
@@
In target-aware federated learning, the goal is to train a model to predict well in a target domain $T(X,Y)$ without access to samples from $T$. Formally, our objective is to minimize the expected target risk, $R_{T}$ of a classifier $h_{\theta}:\mathcal{X}\rightarrow\mathcal{Y}$, with respect to a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$,
$\underset{\theta}{\mbox{minimize}}\;R_{T}(h_{\theta})\coloneqq\underset{(X,Y)\sim T}{\mathbb{E}}\left[\ell(h_{\theta}(X),Y)\right]\leavevmode\nobreak\ .$ (2)
To make solving (2) possible, we assume that target and client label marginal distributions $T(Y),\{S_{i}(Y)\}$ are known to the central server. This is much less restrictive than it sounds: (i) Estimating each client label distribution $S_{i}(Y)$ merely involves computing the proportion of each label in the client sample $D_{i}$, (ii) The target client (domain) may have collected label statistics without logging context features $X$. In our retail example, the label distribution corresponds to the proportion of sales $T(Y=y)$ of each product category $y$, and many companies store this information without logging customer features $X$. The central server is given access to all label distributions to facilitate the learning process, but these are not available to the clients. Retailers may be hesitant to share their exact sales statistics $T(Y)$ with competitors but could share this information with a neutral third party (central server) responsible for coordinating the federated learning process. There is a privacy-accuracy trade-off in all FL settings. In our experiments, we show that substantial performance improvements can be gained at the small privacy cost of sharing the label marginals with the central server.
@@
As in standard FL, clients $i\neq j$ do not communicate directly with each other directly but interact with the central server through model parameters. While it is technically possible for the server to infer each client’s label distribution $S_{i}(Y)$ based on their parameter updates (Ramakrishna and Dán, 2022), doing so would likely be considered a breach of trust in practical applications and sharing would be preferred.
@@
We assume that the distributional shifts between clients and the target are restricted to label shift—while the label distributions vary across clients and the target, the class-conditional input distributions are identical.
@@
For the client distributions $S_{1},...,S_{M}$ and the target distribution $T$,
@@
This setting has been well studied in non-federated learning, see e.g., (Lipton et al., 2018). In the retail example, label shift means that the proportion of sales across product categories ($S_{i}(Y)$ and $T(Y)$) varies between different retailers and the target, but that the pattern of customers who purchase items in each category ($S_{i}(X\mid Y)$ and $T(X\mid Y)$) remain consistent. In other words, although retailers may sell different quantities of products across categories, the characteristics of customers buying a particular product (conditional on the product category) are assumed to be the same. Note that both label shift and covariate shift may hold, that is, there are cases where $\forall i:S_{i}(X\mid Y)=T(X\mid Y)$ and $S_{i}(Y\mid X)=T(Y\mid X)$, but $S_{i}(X),T(X)$ differ, such as when the labeling function is deterministic.
2.1 Limitations of classical aggregation
When either all clients $\{S_{i}\}$ or their aggregate $\bar{S}$, see (1), are identical in distribution to the target domain, the empirical risk on aggregated client data is identical in distribution ($\overset{d}{=}$) to the empirical risk of a hypothetical data set $D_{T}=\{(x_{T,j},y_{T,j})\}_{j=1}^{n_{T}}$ drawn from the target domain,
$\hat{R}\coloneqq\sum_{i=1}^{M}\sum_{j=1}^{n_{i}}\frac{\ell\left(h(x_{i,j}),y_{i,j}\right)}{N}\overset{d}{=}\sum_{j=1}^{n_{T}}\frac{\ell\left(h(x_{T,j}),y_{T,j}\right)}{n_{T}}\eqqcolon\hat{R}_{T}$
Thus, if each client performs a single gradient descent update, the mean of these, weighted by the client sample sizes, is equal in distribution to a centralized batch update for the target domain, given the previous parameter value. This property justifies the federated stochastic gradient (FedSGD) and federated averaging principles (McMahan et al., 2017), both of which aggregate parameter updates in this way,
@@ 2.1 Limitations of classical aggregation
Unfortunately, when the target domain $T$ is not the aggregate of clients $\bar{S}$, the aggregate risk gradient $\nabla\hat{R}$ and, therefore, the FedSGD update are no longer unbiased gradients and updates for the risk in the target domain. As we see in Table 1, this can have large effects on model quality.
Our central question is: How can we aggregate the parameter updates $\theta_{i,t}$ of the $M$ clients, whose data sets are drawn from distributions $S_{1},...,S_{M}$, such that the resulting federated learning algorithm minimizes the target risk, $R_{T}$?
3 FedPALS: Target-aware adjustment for label shift
Figure 1: Illustration of the target label marginal $T$ and client marginals $S_{1},...,S_{4}$ in a ternary classification task, $Y\in\{0,1,2\}$. A: there are fewer clients than labels, $M<K$, and $T\not\in\mathrm{Conv}(S)$; $\alpha^{0}S$ is a projection of $T$ onto $\mathrm{Conv}(S)$. B: $T\in\mathrm{Conv}(S)$ and coincides with $\alpha^{0}S$. In both cases, the label marginal $\alpha^{FA}S$ implied by FedAvg is further from the target distribution.
Next, we develop a model aggregation strategy for target-aware federated learning. Under Assumption 1 (label shift), the target risk is a weighted sum of class-conditional client risks,
$\forall i:R_{T}(h)$ $=\sum_{y=1}^{K}T(y)\mathbb{E}_{S_{i}}[\ell(h(X),y)\mid Y=y]\leavevmode\nobreak\ .$
In centralized learning, this insight is often used to re-weight the training objective in a source domain $S(y)$ by the importance ratio $T(y)/S(y)$ (Lipton et al., 2018; Japkowicz and Stephen, 2002). That is not an option here since $T(Y)$ is not revealed to the clients. For now, assume instead that the target label distribution is covered by the convex hull of the set of client label distributions $S=\{S_{i}(Y)\}_{i=1}^{M}$.
@@ 3 FedPALS: Target-aware adjustment for label shift
The target label distribution $T(Y)$ is covered by the convex hull of client label distributions $S_{1}(Y),...,S_{M}(Y)$, that is $T\in\mathrm{Conv}(S)$, or
$\exists\alpha^{c}\in\Delta^{M-1}:T(y)=\sum_{i=1}^{M}\alpha_{i}^{c}S_{i}(y)\quad\forall y\in[K]\leavevmode\nobreak\ .$ (5)
Under label shift, Assumption 2 implies that $T(X,Y)=\sum_{i=1}^{M}\alpha^{c}_{i}S_{i}(X,Y)$, as well. Thus, under Assumptions 12, we have for any $\alpha^{c}$ satisfying (5),
$R_{T}(h)$ $=\sum_{y=1}^{K}\left(\sum_{i=1}^{M}\alpha^{c}_{i}S_{i}(y)\right)\mathbb{E}[\ell(h(X),y)\mid Y=y]$ (6)
$=\sum_{i=1}^{M}\alpha^{c}_{i}R_{S_{i}}(h)\leavevmode\nobreak\ .$ (7)
Consequently, aggregating client updates with weights $\alpha^{c}$ will be an unbiased estimate of the update.
@@
Consider a single round $t$ of federated learning in the batch stochastic gradient setting with learning rate $\eta$. Each client $i\in[M]$ is given parameters $\theta_{t}$ by the server, computes their local gradient, and returns the update $\theta_{i,t}=\theta_{t}-\eta\nabla_{\theta}\hat{R}_{i}(h_{\theta_{t}})$. Let Assumptions 12 hold and $\alpha^{c}$ satisfy (5). Then, the aggregate update $\theta_{t+1}=\sum_{i=1}^{M}\alpha^{c}_{i}\theta_{i,t}$ satisfies
@@
where $\theta_{t+1}^{T}=\theta_{t}-\eta\nabla_{\theta}\hat{R}_{T}(h_{\theta_{t}})$ is the batch stochastic gradient descent (SGD) update for $\hat{R}_{T}$ that would be obtained with a sample from the target domain. A proof is in Appendix C.
By Proposition 1, we may compute unbiased parameter updates for the target domain by replacing the aggregation step of FedSGD with aggregation weighted according to $\alpha^{c}$. In practice, many federated learning systems, including FedAvg, allow clients several steps of local optimization (e.g., an epoch) before aggregating the parameter updates at the server. Strictly speaking, this is not justified by Proposition 1, but we find in all experiments that aggregating client updates computed over an epoch performs very well, see Section 5.
In applications, the target may not be covered by clients (Assumption 2 may not hold), and $\alpha^{c}$ may not exist. For example, if the target label marginal $T(y)$ is sparse, only clients with exactly the same sparsity pattern as $T$ can be used in a convex combination $\alpha^{c}S=T$. That is, if we aim to classify images of animals and $T$ contains no tigers, then no clients contributing to the combination can have data containing tigers. Since $\{S_{i}(Y)\}_{i=1}^{M},T(Y)$ are known to the server, it is straightforward to verify Assumption 2.
A pragmatic choice when Assumption 2 is violated is to look for the convex combination $\alpha^{0}$ that most closely aligns with the target label distribution, and use that for aggregation,
$\alpha^{0}=\operatorname*{arg\,min}_{\alpha\in\Delta^{M-1}}\bigg{\|}\sum_{i=1}^{M}\alpha_{i}S_{i}(Y)-T(Y)\bigg{\|}_{2}^{2}\quad$ (8)
@@
In light of the above, we propose a client aggregation step such that the weighted sum of clients’ label distributions will a) closely align with the target label distribution, and b) minimize the variance due to weighting using the inverse of the ESS. For a given regularization parameter $\lambda\in[0,\infty)$, we define weights $\alpha^{\lambda}$ as the solution to the following problem
@@
with aggregate client parameters as $\theta^{\lambda}_{t+1}=\sum_{i=1}^{M}\alpha^{\lambda}_{i}\theta_{i,t}$. We refer to this strategy as Federated learning with Parameter Aggregation for Label Shift (FedPALS).